OpenStudy (anonymous):

x=3, r=3 a=? a/3 (xr2)=6x

jhonyy9 (jhonyy9):

a -- (3*9)=6*3 if 2 next r is exponent 3 27a=54 a=2

OpenStudy (anonymous):

a/3 (3*3^2) = 6*3 a/3 *27 = 18 a *27 = 54 a = 2

jhonyy9 (jhonyy9):

can you understanding it ?